Circular No. 8391

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 K4 (LINEAR) M. L. Sitko, University of Cincinnati; R. W. Russell, D. K. Lynch,

and D. L. Kim, Aerospace Corporation; and R. B. Perry, Langley Research Center, NASA, report on further infrared spectrophotometry of comet C/2003 K4 (cf. IAUC 8361, 8378), obtained on Aug. 5.3 and 6.3 UT using the Aerospace Corporation’s BASS spectrograph at the Infrared Telescope Facility 3-m reflector. A smooth continuum was seen between 8 and 13 μm, with a possible weak silicate emission band superimposed. Underlying blackbodies with temperatures of 250 ± 5 K and 245 ± 5 K were fitted to the continuum fluxes at 8.4 and 12 μm on Aug. 5 and 6, respectively. These grain temperatures are ∼ 8–10 percent higher than that of an equilibrium blackbody at the comet’s heliocentric distance. Using the 10.2–10.7-μm region to calculate a silicate feature-to-continuum ratio, this ratio was 1.09 ± 0.03 on Aug. 5, and 1.03 ± 0.03 on Aug. 6. On Aug. 5, a weak feature due to crystalline olivine may have been present at 11.2 μm. Scattered solar radiation was evident at wavelengths shorter than 4 μm. The observed magnitudes determined between 10.2 and 10.7 μm, using the instrument’s 3′′.4 aperture and 18′′ nod, were 2.8 and 2.7 on Aug. 5 and 6, respectively (± 0.03 mag, the errors being dominated by the presence of variable sky transparency during the observations).

SUPERNOVA 2004dj IN NGC 2403 A. V. Filippenko and W. Li, University of California, Berkeley; P.

Challis, Harvard-Smithsonian Center for Astrophysics; and S. D. Van Dyk, Infrared Processing and Analysis Center, California Institute of Technology, report: “We have examined optical images of SN 2004dj obtained on Aug. 17 UT with the Wide Field Camera of the Advanced Camera for Surveys aboard the Hubble Space Telescope. Although the supernova is very saturated in these images, it is clearly located within the compact star cluster Sandage 96, confirming the suggestion made on IAUC 8385 and justified in detail by Maíz-Apellániz et al. (http://xxx.lanl.gov/abs/astroph/?0408265). The tentative identification of a different progenitor star made by Li et al. (IAUC 8388) was based on the radio position reported on IAUC 8379 by Stockdale et al., which is possibly affected by inclusion of extended radio emission in the host galaxy (observed in the VLA-D configuration at 8.46 GHz).”

2004 August 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green